# Low-Latency and Low-Complexity MLSE for Short-Reach Optical Interconnects

### v1

@@ Abstract
To meet the high-speed, low-latency, and low-complexity demand for optical interconnects, simplified layered 2-step maximum likelihood sequence estimation (L2S-MLSE) is proposed in this paper. Simplified L2S-MLSE combines computational simplification and reduced state in L2S-MLSE. L2S-MLSE with a parallel sliding block architecture reduces latency from linear order to logarithmic order. Computational simplification reduces the number of multipliers from exponential order to linear order. Incorporating the reduced state with computational simplification further decreases the number of adders and comparators. The simplified L2S-MLSE is evaluated in a 112-Gbit/s PAM4 transmission over 2-km standard single-mode fiber. Experimental results show that the simplified L2S-MLSE significantly outperforms the FFE-only case in bit error ratio (BER) performance. Compared with simplified 1-step MLSE, the latency of simplified L2S-MLSE is reduced from 34 delay units in linear order to 7 delay units in logarithmic order. The simplified scheme in L2S-MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order without BER performance deterioration, while reducing the number of adders and comparators to 37.2% and 8.4%, respectively, with nearly identical BER performance.
@@ I Introduction
In the DD-FTN scheme, MLSE not only eliminates the residual ISI from FFE, but also eliminates the ISI introduced by the post-filter. However, the performance improvement achieved by MLSE comes with high computational complexity of the Viterbi algorithm. The research about reducing the computational complexity of MLSE is mainly divided into two categories: reducing the number of states and reducing the number of multipliers. To reduce the number of states, the output of FFE can be used as a pre-decision value to reduce the number of candidate states in MLSE [hoon, jing, Hiroki]. Through the pre-decision value, the candidate states with high probability are reserved, while the other states are discarded to reduce the trellis size. Moreover, the decision region whether the reduced-state MLSE is applied can be iteratively determined by error rate [li]. To reduce the number of multipliers, in the branch metric (BM) calculation of MLSE, the lookup table is used to replace the convolution calculation [dai]. The combination of lookup table and reduced state can further reduce the computational complexity [jing2, li2]. The piecewise linear formula or absolute value can also be used to replace the squaring operation in BM calculation [dai, cai].
@@
Besides the complexity, during the add-compare-select (ACS) calculation of MLSE, the current BM should continuously add the previous accumulated metrics in series. This serial calculation structure leads to the continuous increase of latency. With the rapid development of latency-constrained services, such as augmented reality, virtual reality, autonomous driving, and real-time gaming, the demand for low-latency data center optical interconnects is becoming urgent [latency1, latency2, latency3]. Fettweis and Meyr first proposed the look-ahead M-step technique, which can effectively break the latency bottleneck of ACS operation in the conventional Viterbi algorithm [M1, M2, M3]. Inspired by the look-ahead M-step technique, the layered look-ahead Viterbi decoding architecture is used to reduce the latency [M4, M5]. Although the look-ahead M-step technique efficiently decreases the latency, its computational complexity still remains difficult to accept, especially when the modulation order of the signal, the number of channel coefficients in the post-filter, and the number of steps in the M-step are increased.
@@
To meet the high-speed, low-complexity, and low-latency demand for optical interconnects, the simplified layered 2-step maximum likelihood sequence estimation (L2S-MLSE) is proposed in this paper. The latency advantage of the layered 2-step structure and the complexity reduction of simplified L2S-MLSE are analyzed in detail. The bit error ratio (BER) performance of simplified L2S-MLSE is experimentally verified in a 112-Gb/s PAM4 transmission over 2-km SSMF. The main contributions of this paper are as follows:
@@
  • The proposed simplified L2S-MLSE combines the computational simplification and the reduced state in L2S-MLSE. The latency is reduced from linear to logarithmic order through L2S-MLSE with a parallel sliding block architecture. Computational simplification reduces the number of multipliers from exponential to linear order. Incorporating the reduced state further decreases the number of adders and comparators.
@@
  • Experimental results show that compared with simplified 1-step MLSE (1S-MLSE), the latency of simplified L2S-MLSE is decreased from 34 delay units in linear order to 7 delay units in logarithmic order without performance deterioration. Compared with L2S-MLSE, the simplified L2S-MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order, while the number of adders and comparators is reduced to 37.2% and 8.4%, respectively.
@@
The rest of this paper is organized as follows. The principle of parallel 1S-MLSE is introduced in Section II. The parallel L2S-MLSE with low-latency property is described in Section III. In Section IV, the proposed simplified L2S-MLSE is described. In Section V, we present the experimental setup of 112-Gb/s PAM4 transmission over 2-km SSMF. In Section VI, we analyze the parameter settings, BER performance, latency and computational complexity of simplified L2S-MLSE in the experiment. Finally, the paper is concluded in Section VII.
Figure 1: (a) Block diagram of three steps in MLSE. (b) Trellis diagram of PAM4 signal with 2-tap channel response. Sel.: Select.
$PM_{n}(s_{n})=\min\{$ $PM_{n-1}(-3)+BM_{n}(-3,s_{n}),PM_{n-1}(-1)+BM_{n}(-1,s_{n}),$ (3)
$PM_{n-1}(1)\hskip 8.0pt+BM_{n}(1,s_{n}),\hskip 8.0ptPM_{n-1}(3)\hskip 8.0pt+BM_{n}(3,s_{n})\}$
@@ II-A Signal Processing Procedure in MLSE
MLSE employs the Viterbi algorithm to search for the most-likely state transition sequence in the state trellis [viterbi1, viterbi2], which is usually placed after the FFE and post-filter [zhong]. The FFE can compensate for ISI at the cost of noise enhancement. The enhanced noise and the equalized signal are suppressed by the post-filter. Meanwhile, the post-filter can approximate the original channel response with the known shortened coefficients. Then, the signal filtered by the post-filter is fed into MLSE with the known ISI coefficients. The signal processing procedure of MLSE includes three steps, as shown in Fig. 1(a). The three steps are BM calculation, ACS, and survivor path selection.
Figure 2: Sliding block architecture for (a) serial MLSE and (b) parallel MLSE. R: Received symbols; O: Overlapped symbols; D: Data symbols.
1) BM Calculation: In the BM calculation, the BM at time index $n$ can be calculated as
$BM_{n}=\left|y_{n}-\sum_{k=0}^{L-1}{\alpha_{k}s_{n-k}}\right|^{2}$ (1)
where $s_{n}$ and $y_{n}$ are the transmitted signal and received signal at time index $n$. The vector $\bm{\alpha}$ represents the coefficients of the channel response, which includes the elements from $\alpha_{0}$ to $\alpha_{L-1}$, where $\alpha_{0}$ is equal to 1. For the PAM-$M$ signal, each transmitted signal $x_{n}$ at time index $n$ has $M$ possible different states. With $L$ coefficients of the channel response, the trellis has $M^{L-1}$ different states at each time index, and each previous state at time index n1n-1 extends $M$ different branches to the current state at time index $n$. Thus, the total number of BMs from time index n1n-1 to time index $n$ is $M^{L}$. To reduce computational complexity, a 2-tap post-filter ($L=2$) is often used, as it minimizes the number of BMs. For PAM4 modulation with a 2-tap channel response, the BM is calculated as
$BM_{n}(s_{n-1},s_{n})=\left|y_{n}-\left(\alpha s_{n-1}+s_{n}\right)\right|^{2}$ (2)
where $s_{n-1}\in\{-3,-1,1,3\}$ and $s_{n}\in\{-3,-1,1,3\}$. $BM_{n}(s_{n-1},s_{n})$ represents the BM from state $s_{n-1}$ to state $s_{n}$ at time index $n$. Eq. (2) totally includes 16 variable multipliers, 1 constant multiplier for $3\times\alpha$, and 32 adders.
2) ACS Operation: The objective of ACS is to find the best path at each state. Firstly, the accumulated value of previous BMs, which is named as path metric (PM) for each state, is added with the current BM extended from the corresponding state. Secondly, as shown in Fig. 1(b), four different paths converge to the same state, and the path with the minimal PM is chosen at each state (an example is illustrated as a solid line). In conclusion, the ACS computing process can be expressed as Eq. (3), where $PM_{n}(s_{n})$ represents the PM of state $s_{n}$ at time index $n$. The minimal PM value at each state serves as the previous accumulated value for the next ACS calculation. Each ACS unit, as shown in Eq. (3), includes 4 adders and 3 comparators. In this 4-to-1 selection, the first two comparators are used for two 2-to-1 comparisons. Their results are input into the third comparator to obtain the 4-to-1 result. The signal $s_{n}$ includes 4 states, thus the total 4 ACS units at time index $n$ include 16 adders and 12 comparators.
@@
The received symbols calculated in the successive ACS should have a finite length. In the conventional MLSE, when the receiver receives a frame of symbols, the received symbols are processed by serially segmented sliding blocks. As illustrated in Fig. 2(a), each overlapped sliding block of $N$ symbols includes $R$ received symbols followed by $O$ overlapped symbols [TB]. In each sliding block, when the ACS calculation completes at the last overlapped symbol, the final survivor path with $N=R+O$ symbols is obtained. Then, the $R$ received symbols are decoded into $D$ data symbols. Only the results of data symbols are reserved, and the overlapped symbols are removed. The PM of $D$ data symbols is transmitted to the next sliding block as the initial accumulated BM.
@@ II-B Parallel 1S-MLSE: Analysis of Latency and Complexity
In practical MLSE implementation, the received symbols should be processed in parallel [sliding]. As shown in Fig. 2(b), the decision of the state at time index $n$ only depends on the trellis from nOn-O to $n+O$ [sliding]. With the overlap before and behind the data, the received symbols in different blocks can be processed in parallel. Therefore, the received symbols in different blocks are processed through parallel BM calculation, parallel ACS, and parallel survivor path selection. In our paper, all the MLSE algorithms described in the following are based on the parallel sliding block architecture. For the parallel MLSE described in this section, although the received symbols in different blocks are processed in parallel, the symbols within the same block are still processed one after another serially in ACS. To distinguish from the 2-step processing method described in Section III, the parallel MLSE described in this section is termed 1S-MLSE, as symbols within the same block are still processed one after another by 1-step.
The architecture of 1S-MLSE is shown in Fig. 3. One of the parallel sliding blocks with the symbol length of $N=O+R+O$ is analyzed. The BM of all the $N$ received symbols can be calculated in parallel. The PM of the previous time must be transferred to the next time during the ACS operation (the initial PM for the first ACS is zero). Consequently, the prepared BM must wait sequentially as illustrated in Fig. 3. For the block with $N$ received symbols, the latency of ACS should be $N$ delay units, which increases with the increase of block length. Both the BM calculation and survivor path selection introduce the latency of $1$ delay unit, resulting in the overall latency of $N+2$ delay units for 1S-MLSE.
@@
Figure 3: Signal processing architecture for 1S-MLSE in an NN-symbol sliding block based on serial ACS operation.
III Low-Latency L2S-MLSE
@@
In the 1S-MLSE, symbols in the ACS calculation should be processed one after another. The latency of ACS for $N$ received symbols in a sliding block is $N$ delay units, which is increased with the increase of the value $N$. Inspired by the look-ahead $M$-step technique [M1, M2, M3, M4, M5], L2S-MLSE with parallel sliding block architecture can be applied to reduce the latency. Fig. 4(a) illustrates the processing details of L2S-MLSE with $N$ received symbols. The main difference between 1S-MLSE and L2S-MLSE is in the process of ACS. In the 1S-MLSE, one group of $16$ BMs from time index n1n-1 to time index $n$ is processed by 4 ACS units at a time. In the L2S-MLSE, two successive steps of $32$ BMs from time index n1n-1 to time index $n+1$ are processed by 16 ACS units in the first layer. After the first layer, the output of two groups of accumulated BM is still processed by 16 ACS units in the next layer. Finally, the last two groups of accumulated BM are input into 16 ACS units, and the survivor path is found from the outputs of these 16 ACS units. The L2S-MLSE trellis in Fig. 4(b) illustrates an example with a symbol length $N=8$. Owing to the parallel processing, it is obvious that the latency of ACS is reduced from $N$ delay units to $\log_{2}(N)$ delay units.
Figure 4: (a) Signal processing architecture for L2S-MLSE in an NN-symbol sliding block based on layered parallel ACS operation. BM: branch metric; ACS: add-compare-select. (b) Example of an L2S-MLSE trellis with $N=8$. L1: First layer; L2: Second layer; L3: Third layer.
$BM_{n+1}(s_{n-1},s_{n+1})=\min\{$ $BM_{n}(s_{n-1},-3)+BM_{n+1}(-3,s_{n+1}),BM_{n}(s_{n-1},-1)+BM_{n+1}(-1,s_{n+1}),$ (4)
$BM_{n}(s_{n-1},1)\hskip 8.0pt+BM_{n+1}(1,s_{n+1}),\hskip 8.0ptBM_{n}(s_{n-1},3)\hskip 7.2pt+BM_{n+1}(3,s_{n+1})\}$
@@
1) BM Calculation and ACS of the First Layer: The signal processing procedure of L2S-MLSE still includes BM calculation, ACS, and survivor path selection. The BM calculation of L2S-MLSE is the same as that of 1S-MLSE. The details of the ACS calculation of the first layer are shown in Fig. 5. In the first layer, the BMs from time index n1n-1 to time index $n+1$ that share the same initial and final states are summed together. For the 16 ACS units that share the same initial and final states, the calculation can be expressed as Eq. (4). Since all four paths in Eq. (4) start from the same state, their BMs are comparable regardless of the previous PM. In each ACS unit, 4 adders and 3 comparators are included. In the overall 16 ACS units depicted in Fig. 5, 64 adders and 48 comparators are included. In the first layer with $N$ symbols, the overall groups of 16 ACS units are $N/2$. In conclusion, with symbol length of $N$, the overall computational complexity of the BM calculation and ACS operation in the first layer includes $16N$ variable multipliers, $1$ constant multipliers, $48N+16$ adders, and $24N$ comparators.
@@
2) ACS of the Other Layers and Survivor Path Selection: After the first layer, the accumulated BM are transferred layer by layer. The overall groups of 16 ACS units are decreased from the first layer of $N/2$ to the following layers of $N/4$, $N/8$, and so on. In the last layer, the survivor path is found from the outputs of the last group of 16 ACS units. The survivor path selection requires 15 comparators and introduces the latency of $1$ delay unit. Therefore, the overall latency of L2S-MLSE is $\log_{2}(N)+2$ delay units. The overall computational complexity of L2S-MLSE is $16N$ variable multipliers, 1 constant multiplier, 80N4880N-48 adders, and 48N3348N-33 comparators. Compared with 1S-MLSE, the latency of L2S-MLSE is reduced from linear order to logarithmic order at the cost of increased adders and comparators.
IV Simplified L2S-MLSE

### v2

@@ Abstract
To meet the high-speed, low-latency, and low-complexity demand for optical interconnects, simplified maximum likelihood sequence estimation (MLSE) is proposed in this paper. Simplified MLSE combines computational simplification and reduced state in MLSE. MLSE with a parallel sliding block architecture reduces latency from linear order to logarithmic order. Computational simplification reduces the number of multipliers from exponential order to linear order. Incorporating the reduced state with computational simplification further decreases the number of adders and comparators. The simplified MLSE is evaluated in a 112-Gbit/s PAM4 transmission over 2-km standard single-mode fiber. Experimental results show that the simplified MLSE significantly outperforms the FFE-only case in bit error ratio (BER) performance. Compared with simplified 1-step MLSE, the latency of simplified MLSE is reduced from 34 delay units in linear order to 7 delay units in logarithmic order. The simplified scheme in MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order without BER performance deterioration, while reducing the number of adders and comparators to 37.2% and 8.4%, respectively, with nearly identical BER performance.
@@ I Introduction
In the DD-FTN scheme, MLSE not only eliminates the residual ISI from FFE, but also eliminates the ISI introduced by the post-filter. However, the performance improvement achieved by MLSE comes with high computational complexity of the Viterbi algorithm. The research about reducing the computational complexity of MLSE is mainly divided into two categories: reducing the number of states and reducing the number of multipliers. To reduce the number of states, the output of FFE can be used as a pre-decision value to reduce the number of candidate states in MLSE [26, 28, 22]. Through the pre-decision value, the candidate states with high probability are reserved, while the other states are discarded to reduce the trellis size. Moreover, the decision region whether the reduced-state MLSE is applied can be iteratively determined by error rate [18]. To reduce the number of multipliers, in the branch metric (BM) calculation of MLSE, the lookup table is used to replace the convolution calculation [4]. The combination of lookup table and reduced state can further reduce the computational complexity [31, 17]. The piecewise linear formula or absolute value can also be used to replace the squaring operation in BM calculation [4, 5].
@@
Besides the complexity, during the add-compare-select (ACS) calculation of MLSE, the current BM should continuously add the previous accumulated metrics in series. This serial calculation structure leads to the continuous increase of latency. With the rapid development of latency-constrained services, such as augmented reality, virtual reality, autonomous driving, and real-time gaming, the demand for low-latency data center optical interconnects is becoming urgent [23, 21, 16]. Fettweis and Meyr first proposed the look-ahead M-step technique, which can effectively break the latency bottleneck of ACS operation in the conventional Viterbi algorithm [6, 7, 8]. Inspired by the look-ahead M-step technique, the layered look-ahead Viterbi decoding architecture is used to reduce the latency [13, 19]. Although the look-ahead M-step technique efficiently decreases the latency, its computational complexity still remains difficult to accept, especially when the modulation order of the signal, the number of channel coefficients in the post-filter, and the number of steps in the M-step are increased.
@@
To meet the high-speed, low-complexity, and low-latency demand for optical interconnects, the simplified MLSE is proposed in this paper. The latency advantage of the structure and the complexity reduction of simplified MLSE are analyzed in detail. The bit error ratio (BER) performance of simplified MLSE is experimentally verified in a 112-Gb/s PAM4 transmission over 2-km SSMF. The main contributions of this paper are as follows:
@@
  • The proposed simplified MLSE combines the computational simplification and the reduced state in MLSE. The latency is reduced from linear to logarithmic order through MLSE with a parallel sliding block architecture. Computational simplification reduces the number of multipliers from exponential to linear order. Incorporating the reduced state further decreases the number of adders and comparators.
@@
  • Experimental results show that compared with simplified 1-step MLSE (1S-MLSE), the latency of simplified MLSE is decreased from 34 delay units in linear order to 7 delay units in logarithmic order without performance deterioration. Compared with MLSE, the simplified MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order, while the number of adders and comparators is reduced to 37.2% and 8.4%, respectively.
@@
The rest of this paper is organized as follows. The principle of parallel 1S-MLSE is introduced in Section II. The parallel MLSE with low-latency property is described in Section III. In Section IV, the proposed simplified MLSE is described. In Section V, we present the experimental setup of 112-Gb/s PAM4 transmission over 2-km SSMF. In Section VI, we analyze the parameter settings, BER performance, latency and computational complexity of simplified MLSE in the experiment. Finally, the paper is concluded in Section VII.
@@ II-A Signal Processing Procedure in MLSE
MLSE employs the Viterbi algorithm to search for the most-likely state transition sequence in the state trellis [9, 15], which is usually placed after the FFE and post-filter [29]. The FFE can compensate for ISI at the cost of noise enhancement. The enhanced noise and the equalized signal are suppressed by the post-filter. Meanwhile, the post-filter can approximate the original channel response with the known shortened coefficients. Then, the signal filtered by the post-filter is fed into MLSE with the known ISI coefficients. The signal processing procedure of MLSE includes three steps. The three steps are BM calculation, ACS, and survivor path selection.
1) BM Calculation: In the BM calculation, $s_{n}$ and $y_{n}$ are the transmitted signal and received signal at time index $n$. The vector $\bm{\alpha}$ represents the coefficients of the channel response, which includes the elements from $\alpha_{0}$ to $\alpha_{L-1}$, where $\alpha_{0}$ is equal to 1. For the PAM-$M$ signal, each transmitted signal $x_{n}$ at time index $n$ has $M$ possible different states. With $L$ coefficients of the channel response, the trellis has $M^{L-1}$ different states at each time index, and each previous state at time index n1n-1 extends $M$ different branches to the current state at time index $n$. Thus, the total number of BMs from time index n1n-1 to time index $n$ is $M^{L}$. To reduce computational complexity, a 2-tap post-filter ($L=2$) is often used, as it minimizes the number of BMs.
2) ACS Operation: The objective of ACS is to find the best path at each state. Firstly, the accumulated value of previous BMs, which is named as path metric (PM) for each state, is added with the current BM extended from the corresponding state. Secondly, four different paths converge to the same state, and the path with the minimal PM is chosen at each state (an example is illustrated as a solid line). The minimal PM value at each state serves as the previous accumulated value for the next ACS calculation. Each ACS unit includes 4 adders and 3 comparators. In this 4-to-1 selection, the first two comparators are used for two 2-to-1 comparisons. Their results are input into the third comparator to obtain the 4-to-1 result. The signal $s_{n}$ includes 4 states, thus the total 4 ACS units at time index $n$ include 16 adders and 12 comparators.
@@
The received symbols calculated in the successive ACS should have a finite length. In the conventional MLSE, when the receiver receives a frame of symbols, the received symbols are processed by serially segmented sliding blocks. Each overlapped sliding block of $N$ symbols includes $R$ received symbols followed by $O$ overlapped symbols [2]. In each sliding block, when the ACS calculation completes at the last overlapped symbol, the final survivor path with $N=R+O$ symbols is obtained. Then, the $R$ received symbols are decoded into $D$ data symbols. Only the results of data symbols are reserved, and the overlapped symbols are removed. The PM of $D$ data symbols is transmitted to the next sliding block as the initial accumulated BM.
@@ II-B Parallel 1S-MLSE: Analysis of Latency and Complexity
In practical MLSE implementation, the received symbols should be processed in parallel [3]. The decision of the state at time index $n$ only depends on the trellis from nOn-O to $n+O$ [3]. With the overlap before and behind the data, the received symbols in different blocks can be processed in parallel. Therefore, the received symbols in different blocks are processed through parallel BM calculation, parallel ACS, and parallel survivor path selection. In our paper, all the MLSE algorithms described in the following are based on the parallel sliding block architecture. For the parallel MLSE described in this section, although the received symbols in different blocks are processed in parallel, the symbols within the same block are still processed one after another serially in ACS. To distinguish from the processing method described in Section III, the parallel MLSE described in this section is termed 1S-MLSE, as symbols within the same block are still processed one after another by 1-step.
For the architecture of 1S-MLSE, one of the parallel sliding blocks with the symbol length of $N=O+R+O$ is analyzed. The BM of all the $N$ received symbols can be calculated in parallel. The PM of the previous time must be transferred to the next time during the ACS operation (the initial PM for the first ACS is zero). Consequently, the prepared BM must wait sequentially. For the block with $N$ received symbols, the latency of ACS should be $N$ delay units, which increases with the increase of block length. Both the BM calculation and survivor path selection introduce the latency of $1$ delay unit, resulting in the overall latency of $N+2$ delay units for 1S-MLSE.
@@
III Low-Latency MLSE
@@
In the 1S-MLSE, symbols in the ACS calculation should be processed one after another. The latency of ACS for $N$ received symbols in a sliding block is $N$ delay units, which is increased with the increase of the value $N$. Inspired by the look-ahead $M$-step technique [6, 7, 8, 13, 19], MLSE with parallel sliding block architecture can be applied to reduce the latency. The main difference between 1S-MLSE and MLSE is in the process of ACS. In the 1S-MLSE, one group of $16$ BMs from time index n1n-1 to time index $n$ is processed by 4 ACS units at a time. In the MLSE, two successive steps of $32$ BMs from time index n1n-1 to time index $n+1$ are processed by 16 ACS units in the first layer. After the first layer, the output of two groups of accumulated BM is still processed by 16 ACS units in the next layer. Finally, the last two groups of accumulated BM are input into 16 ACS units, and the survivor path is found from the outputs of these 16 ACS units. Owing to the parallel processing, it is obvious that the latency of ACS is reduced from $N$ delay units to $\log_{2}(N)$ delay units.
@@
1) BM Calculation and ACS of the First Layer: The signal processing procedure of MLSE still includes BM calculation, ACS, and survivor path selection. The BM calculation of MLSE is the same as that of 1S-MLSE. In the first layer, the BMs from time index n1n-1 to time index $n+1$ that share the same initial and final states are summed together. Since all four paths start from the same state, their BMs are comparable regardless of the previous PM. In each ACS unit, 4 adders and 3 comparators are included. In the overall 16 ACS units, 64 adders and 48 comparators are included. In the first layer with $N$ symbols, the overall groups of 16 ACS units are $N/2$. In conclusion, with symbol length of $N$, the overall computational complexity of the BM calculation and ACS operation in the first layer includes $16N$ variable multipliers, $1$ constant multipliers, $48N+16$ adders, and $24N$ comparators.
@@
2) ACS of the Other Layers and Survivor Path Selection: After the first layer, the accumulated BM are transferred layer by layer. The overall groups of 16 ACS units are decreased from the first layer of $N/2$ to the following layers of $N/4$, $N/8$, and so on. In the last layer, the survivor path is found from the outputs of the last group of 16 ACS units. The survivor path selection requires 15 comparators and introduces the latency of $1$ delay unit. Therefore, the overall latency of MLSE is $\log_{2}(N)+2$ delay units. The overall computational complexity of MLSE is $16N$ variable multipliers, 1 constant multiplier, 80N4880N-48 adders, and 48N3348N-33 comparators. Compared with 1S-MLSE, the latency of MLSE is reduced from linear order to logarithmic order at the cost of increased adders and comparators.
IV Simplified MLSE